# Fingerprinting Implementations of Cryptographic Primitives and Protocols that Use Post-Quantum Algorithms

### v3

@@ 8 Conclusion
In this work, we demonstrate the feasibility of distinguishing between post-quantum and classical algorithms based on their implementation within cryptographic libraries, SNARK generating libraries and widely used protocols. We classified and identified library implementations for both post-quantum and classical primitives and SNARKS. This classification is achieved by analyzing the CPU cycle counts and memory footprints associated with these algorithms, utilizing ensemble learning models that deliver high accuracy in differentiation. We extended our analysis to the identification of key exchange algorithms employed within various connections, specifically focusing on post-quantum implementations of protocols such as TLS, SSH, OpenID Connect, QUIC, and VPN. To achieve this, we meticulously analyze both self-generated and publicly available packet captures. By filtering for key exchange packets, we are able to accurately discern the specific key exchange algorithms being utilized. Finally, we integrate our fingerprinting methods with QUARTZ and also identify domains in Tranco which use post qauntum TLS handshakes.
Appendix A Integration of digital signatures in secure protocols
Appendix B Simplified Classical Key Exchange in TLS 1.3
Figure 3: Simplified Classical Key Exchange in TLS 1.3
Appendix C Simplified Hybrid Key Exchange in TLS 1.3
Figure 4: Simplified Hybrid Key Exchange in TLS 1.3

### v4

@@ 8 Conclusion
In this work, we demonstrate the feasibility of distinguishing between post-quantum and classical algorithms based on their implementation within cryptographic libraries, SNARK generating libraries and widely used protocols. We classified and identified library implementations for both post-quantum and classical primitives and SNARKs. The classification is done by analyzing the CPU cycle counts and memory footprints associated with these algorithms, utilizing ensemble learning models that deliver high accuracy in differentiation. We extended our analysis to the identification of key exchange algorithms employed within various connections, specifically focusing on post-quantum implementations of protocols such as TLS, SSH, OpenID Connect, QUIC, and VPN. To achieve this, we meticulously analyze both self-generated and publicly available packet captures. By filtering for key exchange packets, we are able to accurately discern the specific key exchange algorithms being utilized. Finally, we integrate our classification methods with Quartz and also identify domains in Tranco which use post qauntum TLS handshakes.